Coral Language Please to thanks

Write a program that first gets a list of 5 integers from input. Then, get another value from the input, and output all integers less than or equal to that value.

Ex: If the input is 50 60 140 200 75 100, the output is:

50 60 75

For coding simplicity, follow every output value by a space, including the last one. Then, output a newline.

Such functionality is common on sites like Amazon, where a user can filter results.
Answer #1

Thanks for the question. Below is the code you will be needing. Let me know if you have any doubts or if you need anything to change.

If you are satisfied with the solution, please leave a +ve feedback : ) Let me know for any help with any other questions.

Thank You!
===========================================================================

integer array(5) userNums
integer i
integer compareWithNum

for i = 0; i < userNums.size; i = i + 1
userNums[i] = Get next input

compareWithNum = Get next input

for i = 0; i < userNums.size; i = i + 1
if userNums[i]<=compareWithNum
Put userNums[i] to output
Put " " to output

1 integer array(5) userNums 2 integer i 3 integer compareWithNum 4 5 for i = 0; i < user Nums.size; i = i + 1 6 userNums[i] =